### v1

# A Review of Modern Recommender Systems Using Generative Models (Gen-RecSys)

Abstract.
Traditional recommender systems (RS) have used user-item rating histories as their primary data source, with collaborative filtering being one of the principal methods. However, generative models have recently developed abilities to model and sample from complex data distributions, including not only user-item interaction histories but also text, images, and videos – unlocking this rich data for novel recommendation tasks. Through this comprehensive and multi-disciplinary survey, we aim to connect the key advancements in RS using Generative Models (Gen-RecSys), encompassing: a foundational overview of interaction-driven generative models; the application of large language models (LLM) for generative recommendation, retrieval, and conversational recommendation; and the integration of multimodal models for processing and generating image and video content in RS. Our holistic perspective allows us to highlight necessary paradigms for evaluating the impact and harm of Gen-RecSys and identify open challenges. A more up-to-date version of the papers are maintained at: https://github.com/yasdel/LLM-RecSys.
Generative Models, Recommender Systems, GANs, VAEs, LLMs, Personalization, Ethical and Societal Considerations
*This work does not relate to the author’s position at Amazon.
1. Introduction
Advancements in generative models have significantly impacted the evolution of recommender systems (RS). Traditional RS, which relied on capturing user preferences and item features within a specific domain — often referred to as “narrow experts” – are now being complemented and, in some instances, surpassed by generative models. These models have introduced innovative ways of conceptualizing and implementing recommendations. Specifically, modern generative models learn to represent and sample from complex data distributions, including not only user-item interaction histories but also text and image content, unlocking these data modalities for novel and interactive recommendation tasks.
Moreover, advances in natural language processing (NLP) through the introduction of large language models (LLMs) such as ChatGPT (OpenAI, 2022) and Gemini (Team et al., 2023) have showcased remarkable emergent capabilities (Wei et al., 2022), including reasoning, in-context few-shot learning, and access to extensive open-world information within their pre-trained parameters. Because of their broad generalist abilities, these pretrained generative models have opened up an exciting new research space for a wide variety of recommendation applications (see Table 1), e.g., enhanced personalization, improved conversational interfaces, and richer explanation generation, among others.
Figure 1. Overview of the areas of interest in generative models in recommendation.
The core of generative models lies in their ability to model and sample from the data distribution they are trained on for various inferential purposes. This enables two primary modes of application for RS:
  1. (1)
    Directly trained models. This approach trains generative models, such as VAE-CF(Variational AutoEncoders for Collaborative Filtering) (Liang et al., 2018), directly on user-item interaction data to predict user preferences, without using large, diverse pretraining datasets. These models learn the probability distribution of items a user might like based on their previous interactions.
  2. (2)
    Pretrained models. This strategy uses models pretrained on diverse data (text, images, videos) to understand complex patterns, relationships, and contexts. Among other methods, this survey covers the use of pretrained generative models for RS in the following settings:
    • Zero- and Few-shot Prompting (cf. Section 3.2.1), with the latter using in-context learning (ICL) for broad understanding without extra training.
    • Fine-Tuning (cf. Section 3.3), adjusting models with specific datasets for tailored recommendations.
    • Retrieval-Augmented Generation (RAG) (cf. Section 3.3), integrating information retrieval with generative modeling for contextually relevant outputs.
    • Embedding for Downstream Training (cf. Section 3.1), generating embeddings or token sequences for complex content representation.
    • Multimodal Approaches (cf. Section 4), using various data types to improve the accuracy and relevance of model recommendation.
Table 1. Example applications of Gen-RecSys methods.
Description Relevant Section
Leveraging text in RS, including descriptions of items and user preferences as well as user-system interactions such as reviews, queries or conversation histories, with examples including generative recommendation and explanation, and conversational recommendation. Sections 3.2, 3.5
Using and generating images for recommendation, reasoning, and content creation. Section 4, Sections 4.4
Applying the emergent reasoning abilities of pre-trained models to recommendation tasks, including in-context learning and tool-augmented reasoning. Sections 3.3, 3.5
Integrating RS with external knowledge sources through retrieval augmented generation. Section 3.3
Selecting or generating informative user-item interactions to improve recommendation model training. Sections 2.3, 2.4
Generating recommendation results with complex structures such as list-wise or page-wise outputs. Sections 2.1, 2.2, 2.3, 2.5
Facilitating conversational recommendation through full NL dialogue. Section 3.5
1.1. Recent Surveys and Our Contributions
Recent Relevant Surveys
Recent surveys have marked significant advancements in the field. We highlight our contributions and distinguish our survey by its comprehensive and unique approach.
  • Deldjoo et al. (2021) explore GAN-based RS across four different recommendation scenarios (graph-based, collaborative, hybrid, context-aware).
  • Li et al. (2023b) explore training strategies and learning objectives of LLMs for RS.
  • Wu et al. (2023b) discuss both the use of LLMs to generate RS input tokens or embeddings as well as the use of LLMs as an RS;
  • Lin et al. (2023) focus on adapting LLMs in RS, detailing various tasks and applications.
  • Fan et al. (2023) overview LLMs in RS, emphasizing pre-training, fine-tuning, and prompting.
  • Vats et al. (2024) review LLM-based RS, introducing a heuristic taxonomy for categorization.
  • Huang et al. (2024a), explore using foundation models (FMs) in RS.
  • Wang et al. (2023b) introduce GeneRec, a next-gen RS that personalizes content through AI generators and interprets user instructions to gather user preferences.
While the mentioned surveys offer crucial insights, their scope is often limited to LLMs (Li et al., 2023b; Wu et al., 2023b; Lin et al., 2023; Fan et al., 2023; Vats et al., 2024) or, more broadly, FMs (Huang et al., 2024a) and/or specific sets of models such as GANs (Deldjoo et al., 2021), without encompassing the wider spectrum of generative models and data modalities. In contrast, the work by (Wang et al., 2023b) provides a more comprehensive survey focused on these broader aspects of Gen-RecSys although their work is mostly on personalized content generation.
Core Contributions.
Figure 1 depicts the organization of our survey on Gen-RecSys. It is categorized into layers that explores data sources, recommendation models, and scenarios with a vertical hierarchy that extends to the evaluation of these systems and the challenges they face. We thus offer a systematic approach to breaking down the recommendation process of Gen-RecSys research into discrete components and methodologies. Our contributions can be briefly summarized as follows.
  1. (1)
    Our survey is broader in scope than the surveys mentioned above, encompassing not just LLMs but a wide array of generative models in RS.
  2. (2)
    We have chosen to classify these models based on the type of data and modality they are used for, such as user-item data (cf. Section 2), text-driven (cf. Section 3), and multimodal (cf. Section 4) models, as shown in the Rec. Scenario layer.
  3. (3)
    Within each modality discussion, we provide an in-depth exploration of deep generative model paradigms as shown in the Model layer, yet with a broader scope that spans multiple contexts and use cases, offering a critical analysis of their roles and effectiveness in respective sections.
  4. (4)
    We study the evaluation of Gen-RecSys (cf. Section 5) with finer details, shedding light on multiple aspects such as benchmarks, evaluation for impact and harm relative to multiple stakeholders, and conversational evaluation. This evaluation framework is particularly notable as it helps to understand the complex challenges intrinsic to Gen-RecSys.
  5. (5)
    We discuss several open research challenges and issues (cf. Section 6). Our survey benefits from the expertise of distinguished scholars and industry practitioners from a variety of institutions and disciplines.
2. Generative Models for Interaction-
Driven Recommendation
Interaction-driven recommendation is a setup where only the user-item interactions (e.g., “user A clicks item B”) are available, which is the most general setup studied in RS. In this setup, we concentrate on the inputs of user-item interactions and outputs of item-recommended lists or grids rather than richer inputs or outputs from other modalities such as recommendation explanations. Even though no textual or visual information is involved, Deep Generative Models (DGMs) (Goodfellow et al., 2014; Kingma and Welling, 2013; Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017; Sohl-Dickstein et al., 2015) still show their unique usefulness. For example, DGMs can augment user-item interactions for model training (Cai and Wang, 2018), use denoising (Wang et al., 2023d) for recommendation, and learn the distributions of recommendation layouts (Xiao et al., 2022).
In this section, we examine the paradigms of DGMs for recommendation tasks with data on user-item interactions, including auto-encoding models (Kingma and Welling, 2013), auto-regressive models (Vaswani et al., 2017; Hochreiter and Schmidhuber, 1997), generative adversarial networks (Goodfellow et al., 2014), diffusion models (Sohl-Dickstein et al., 2015) and more [^1] When discussing DGMs, confusion may arise due to the conflation of model paradigms and architectures. Model architectures, such as multilayer perceptron (Hornik et al., 1989), self-attentive models (Vaswani et al., 2017) can be used across various paradigms (Kang and McAuley, 2018; Liu et al., 2023a)..
2.1. Auto-Encoding Models
Auto-encoding models learn to reconstruct their inputs. This capability allows them to be used for various purposes, including denoising, representation learning, and generation tasks.
2.1.1. Preliminaries: Denoising Auto-Encoding Models
Denoising Auto-Encoding models are a group of models that learn to recover the original inputs from a corrupted version of the inputs. Traditionally, denoising auto-encoding models refer to a group of Denoising Autoencoders (Vincent et al., 2008; Sedhain et al., 2015) with hidden layers as a “bottleneck”. For example, AutoRec (Sedhain et al., 2015) tries to reconstruct the input vector, which is partially observed. More broadly, BERT-like models (Devlin et al., 2019a; Sun et al., 2019; Wu et al., 2016) are also treated as denoising auto-encoding models. Such models recover corrupted (i.e., masked) inputs through stacked self-attention blocks (Sun et al., 2019; He et al., 2021). For example, BERT4Rec (Sun et al., 2019) is trained to predict masked items in given user historical interaction sequences. Therefore, BERT-like (Devlin et al., 2019a) models can be used for next-item prediction in the inference phase (Sun et al., 2019; He et al., 2021).
2.1.2. Variational Auto-Encoding Models
Variational Autoencoders (VAEs) are models that learn stochastic mappings from an input $x$ from a often complicated probability distribution $p$ to a probability distribution $q$. This distribution, $q$, is typically simple (e.g., a normal distribution), enabling the use of a decoder to generate outputs $\hat{x}$ by sampling from $q$ (Kingma and Welling, 2013). VAEs find wide application in traditional RS, particularly for collaborative filtering (Liang et al., 2018), sequential recommendation (Sachdeva et al., 2019) and slate generation (Jiang et al., 2018; Liu et al., 2021; Deffayet et al., 2023). Compared to Denoising Autoencoders (DAEs), VAEs often demonstrate superior performance in collaborative filtering due to stronger modeling assumptions, such as VAE-CF (Liang et al., 2018). Additionally, Conditional VAE (CVAE) (Sohn et al., 2015) models learn distributions of preferred recommendation lists for a given user. This makes them useful for generating those lists beyond a greedy ranking schema. Examples like ListCVAE (Jiang et al., 2018) and PivotCVAE (Liu et al., 2021) use VAEs to generate entire recommendation lists rather than solely ranking individual items.
2.2. Auto-Regressive Models
Given an input sequence $\mathbf{x}$, at step $i$, auto-regressive models (Bengio et al., 2000) learn the conditional probability distribution $p(x_{i}|\mathbf{x}_{<i})$, where $\mathbf{x}_{<i}$ represents the subsequence before step $i$. Auto-regressive models are primarily used for sequence modeling (Bengio et al., 2000; Vaswani et al., 2017; Dieleman et al., 2016). In RS, they find wide applications in session-based or sequential recommendations (Hidasi et al., 2016; Kang and McAuley, 2018), model attacking (Yue et al., 2021), and bundle recommendations (Bai et al., 2019; Hu and He, 2019), with recurrent neural networks (Hidasi et al., 2016; Bai et al., 2019; Hu and He, 2019), self-attentive models (Kang and McAuley, 2018), and more.
2.2.1. Recurrent Auto-Regressive Models
Recurrent neural networks (RNNs) (Hochreiter and Schmidhuber, 1997; Chung et al., 2014) have been use to predict the next item in session-based and sequential recommendations, such as GRU4Rec (Hidasi et al., 2016) and its variants (Hidasi and Karatzoglou, 2018; Yue et al., 2024) (e.g., predicting the next set of items in basket or bundle recommendations, such as set2set (Hu and He, 2019) and BGN (Bai et al., 2019)). Moreover, using the auto-regressive generative nature of recurrent networks, researchers extract model-generated user behavior sequences, which are used in the research of model attacking (Yue et al., 2021).
2.2.2. Self-Attentive Auto-Regressive Models
Self-attentive models replace the recurrent unit with self-attention and related modules, inspired by transformers (Vaswani et al., 2017). This group of models can be used in session-based recommendation and sequential recommendation (Kang and McAuley, 2018; Wu et al., 2020; Lin et al., 2020), next-basket or bundle prediction (Yu et al., 2020), and model attacking (Yue et al., 2021), similar to RNNs. Meanwhile, the benefits of self-attentive models are that they handle long-term dependencies better than RNNs and enable parallel training (Vaswani et al., 2017). Additionally, self-attentive models are the de-facto option for pre-trained models (Devlin et al., 2019a) and large language models (Brown et al., 2020; Wei et al., 2022; Bubeck et al., 2023), which is gaining traction in RS. More details about using such language models for recommendations will be discussed in Section 3.
2.3. Generative Adversarial Networks
Generative adversarial networks (GANs) (Mirza and Osindero, 2014; Goodfellow et al., 2014) are composed of two primary components: a generator network and a discriminator network. These networks engage in adversarial training to enhance the performance of both the generator and the discriminator. GANs are used in RS for multiple purposes (Wang et al., 2017; Cai and Wang, 2018; Chen et al., 2019a). In the interaction-driven setup, GANs are proposed for selecting informative training samples (Cai and Wang, 2018; Wang et al., 2017), for example, in IRGAN (Wang et al., 2017, 2018), the generative retrieval model is leveraged to sample negative items. Meanwhile, GANs synthesize user preferences or interactions to augment training data (Chae et al., 2018; Wang et al., 2019). Additionally, GANs have shown effectiveness in generating recommendation lists or pages, such as (Chen et al., 2019a) in whole-page recommendation settings.
2.4. Diffusion Models
Diffusion models  (Sohl-Dickstein et al., 2015) generate outputs through a two-step process: (1) corrupting inputs into noise via a forward process, and (2) learning to recover the original inputs from the noise iteratively in a reverse process. Their impressive generative capabilities have attracted growing interest from the recommender systems community.
First, a group of works (Wang et al., 2023d; Walker et al., 2022) aims to learn users’ future interaction probabilities through diffusion models. For example, DiffRec (Wang et al., 2023d) predicts users’ future interactions using corrupted noises from the users’ historical interactions. Second, another group of works (Liu et al., 2023d; Wu et al., 2023a) focuses on using diffusion models to augment training sequences, showing promising results in alleviating the data sparsity and long-tail user problems in sequential recommendation.
2.5. Other Generative Models
In addition to the previously mentioned generative models, RS also draw upon other types of generative models. For instance, VASER (Zhong et al., 2020) leverages normalizing flows (Rezende and Mohamed, 2015) (and VAEs (Kingma and Welling, 2013)) for session-based recommendation. GFN4Rec (Liu et al., 2023a), on the other hand, adapts generative flow networks (Bengio et al., 2021; Pan et al., 2023) for listwise recommendation. Furthermore, IDNP (Du et al., 2023) utilizes generative neural processes (Garnelo et al., 2018b, a) for sequential recommendation. In summary, various generative models are explored in RS, even in settings without textual or visual modalities.
3. Large Language Models in Recommendation
While language has been leveraged by content-based RS for over three decades (Lops et al., 2011), the advent of pretrained LLMs and their emergent abilities for generalized, multi-task natural language (NL) reasoning (Brown et al., 2020; Wei et al., 2022; Bubeck et al., 2023) has marked the beginning of a new stage of language-based recommendation. Exploiting textual content is a compelling prospect, since NL constitutes a unified, expressive, and interpretable medium that can represent not only item features or user preferences, but also user-system interactions, recommendation tasks, and external knowledge. Specifically, real items are often associated with rich text content including titles, descriptions, semi-structured textual metadata, and reviews. Similarly, user preferences can be easily articulated in NL in both traditional systems and the emerging paradigm of conversational RS: in the former by leveraging NL components such as reviews, search queries, or liked item descriptions, and in the latter through direct NL user utterances. Pretrained LLMs provide new ways to exploit this textual data, since, in many domains, LLMs have learned useful reasoning abilities for making and explaining item recommendations based on user preferences (Sanner et al., 2023; Sileo et al., 2022; Geng et al., 2022; He et al., 2023). These pretrained abilities can be further augmented through prompting (Sanner et al., 2023; Sileo et al., 2022; Liu et al., 2023c), fine-tuning (Geng et al., 2022; Bao et al., 2023; Harte et al., 2023; Mao et al., 2023; Kang et al., 2023; Cui et al., 2022; Zhang et al., 2023b), retrieval (Hou et al., 2023; Chen, 2023; Wang and Lim, 2023; Dai et al., 2023), and external tools (Wang et al., 2023a; Friedman et al., 2023; Gao et al., 2023).
We next explore key evolving approaches using LLMs in RS in both single- and multi-stage settings, covering:
  • Encoder-only LLM Recommendation: Section 3.1 discusses LLMs used either as highly-scalable dense retrievers or to cross-encode preference and item information for top-K recommendation and rating prediction.
  • LLM-based Generative Recommendation Section 3.2 introduces generative recommendation with LLMs in zero-shot (ZS) and few-shot (FS) settings, as well as approaches to internalize recommendation knowledge through fine-tuning, prompt-tuning, and controllable decoding.
  • LLM/RS Hybrids We explore the complimentary use of LLMs and RS in multi-stage pipelines, with Section 3.3 focusing on retrieval-augmented generation and Section 3.4 discussing LLMs as input generators for RS.
  • Conversational Recommendation: Section 3.5 covers recommendation in NL dialogue, including work on tool augmented LLM systems for of dialogue management, recommendation, explanation, and retrieval.
3.1. Encoder-only LLM Recommendation
3.1.1. Recommendation as Dense Retrieval
To use dense retrieval for recommendation, each item’s text content (e.g., title, description, and/or reviews) is treated as a document, and if no explicit query is available from the user, a common approach is to synthesize a query by concatenating the descriptions of a user’s recently liked items (often titles for conciseness (Harte et al., 2023; Penha and Hauff, 2020; Li et al., 2023c)). Dense retrievers (e.g., BERT (Devlin et al., 2019b), TAS-B (Hofstätter et al., 2021), Condenser (Gao and Callan, 2021)) produce a ranked list of documents given a query by evaluating the similarity (e.g., dot product or cosine similarity) between an encoder-only LLM document embedding and query embedding. Dense retrieval is highly scalable (especially with approximate search libraries like FAISS [^2] https://github.com/facebookresearch/faiss) as documents and queries are encoded separately, allowing for dense vector indexing of documents before querying. Recent work explores recommendation as dense retrieval with off-the-shelf retrievers (Penha and Hauff, 2020; Harte et al., 2023), contrastive BERT fine-tuning for review-based retrieval (Abdollah Pour et al., 2023), the use of a second-level encoder fusing the embedding of a user’s recently liked items into a user embeddings before scoring (Wu et al., 2019; Li et al., 2022b) or fine-tuned components of multi-stage systems (Mysore et al., 2023; Li et al., 2023c; Hou et al., 2023).
3.1.2. Recommendation via LLM Item-Preference Fusion
Several works approach rating prediction by jointly embedding NL item and preference descriptions in LLM cross-encoder architectures with an MLP rating prediction head (Zhang et al., 2021; Yao et al., 2022; Wu et al., 2021; Qiu et al., 2021; Zhang and Wang, 2023). Such fusion-in-encoder methods often exhibit strong performance because they allow interaction between user and item representations, but are much more computationally expensive than dense retrieval and thus may be best used for small item sets or as rerankers (Mysore et al., 2023).
3.2. LLM-based Generative Recommendation
In LLM-driven generative recommendation, tasks are expressed as token sequences – called prompts – which form an input to an encoder-decoder LLM (e.g. GPT3 (Brown et al., 2020), T5 (Raffel et al., 2020), PALM (Chowdhery et al., 2022)). The LLM then generates another token sequence to address the task – with example outputs including: a recommended list of item titles/ids (Mao et al., 2023; Harte et al., 2023; Sanner et al., 2023; Sileo et al., 2022), a rating (Bao et al., 2023; Kang et al., 2023), or an explanation (Ni et al., 2019; Li et al., 2020; Hada and Shevade, 2021; Geng et al., 2022; Li et al., 2023a). Generative recommendation with LLMs relies on pretraining on large text corpora to provide knowledge about a wide range of entities, human preferences, and commonsense knowledge that can be used directly for recommendation or leveraged to improve generalization and reduce domain-specific data requirements for fine-tuning or prompting (Wei et al., 2022; Bubeck et al., 2023).
3.2.1. Zero- and Few- Shot Generative Recommendation
Several recent publications (Kang et al., 2023; Sanner et al., 2023; Sileo et al., 2022; Liu et al., 2023c) have evaluated with off-the-shelf LLM generative recommendation, focusing on domains that are prevalent in the LLM pre-training corpus such as movie and book recommendation. Specifically, these methods construct a prompt with a NL description of user preference (often using a sequence of recently liked item titles) and an instruction to recommend the next $k$ item titles (Sanner et al., 2023; Sileo et al., 2022; Liu et al., 2023c) or predict a rating (Kang et al., 2023; Liu et al., 2023c). While, overall, untuned LLMs underperform supervised CF methods trained on sufficient data (Kang et al., 2023; Sileo et al., 2022), they are competitive in near cold-start settings (Sanner et al., 2023; Sileo et al., 2022). Few-shot prompting (or in-context learning), in which a prompt contains examples of input-output pairs, typically outperforms zero-shot prompting (Sanner et al., 2023).
3.2.2. Tuning LLMs for Generative Recommendation
To improve an LLM’s generative recommendation performance and add knowledge to its internal parameters, multiple works focus on fine-tuning (Geng et al., 2022; Bao et al., 2023; Harte et al., 2023; Mao et al., 2023; Kang et al., 2023) and prompt-tuning (Li et al., 2023a; Cui et al., 2022; Zhang et al., 2023b) strategies. Recent works fine-tune LLMs on NL input/output examples constructed from user interaction history and task descriptions for rating prediction (Bao et al., 2023; Kang et al., 2023) and sequential recommendation (Mao et al., 2023; Harte et al., 2023), or in the case of P5 (Geng et al., 2022), both preceding tasks plus top-$k$ recommendation, explanation generation, and review summarization. Other recommendation works study prompt tuning approaches (Li et al., 2023a; Cui et al., 2022; Zhang et al., 2023b), which adjust LLM behaviour by tuning a set of continuous (or soft) prompt vectors as an alternative to tuning LLM weights.
Generative Explanation
A line of recent work focuses on explanation generation, where prompts are typically constructed from a user’s item or review histories while training explanations are extracted from reviews. Techniques include fine-tuning (Geng et al., 2022; Li et al., 2023a), prompt-tuning (Li et al., 2023a, 2020), and controllable decoding (Ni et al., 2019; Hada and Shevade, 2021), where additional predicted parameters such as ratings steer LLM decoding.
3.3. Retrieval Augmented Recommendation
While adding knowledge to an LLM internal memory through tuning often improves performance, this approach requires a large number of LLM parameters to store knowledge, and re-tuning is needed for every system update. An alternative paradigm is retrieval-augmented generation (RAG) (Lewis et al., 2020; Izacard and Grave, 2020; Borgeaud et al., 2022), where output generation is conditioned on information retrieved from an external knowledge source such as a dense retriever (Section 3.1). RAG methods can better facilitate online updates, reduce hallucinations, and generally require fewer LLM parameters since knowledge can be externalized (Mialon et al., 2023; Borgeaud et al., 2022; Izacard and Grave, 2020). RAG has recently begun to be explored for recommendation, with the most common approach being to first use a retriever or RS to construct a candidate item set based on a user query or interaction history, and then prompt an encoder-decoder LLM to rerank the candidate set (Hou et al., 2023; Chen, 2023; Wang and Lim, 2023; Dai et al., 2023).
3.4. LLM-based Input Generation
Conversely to how RS or retrievers are used in RAG to obtain inputs for LLMs (Section 3.3), LLMs can also be used to generate inputs for RS (Harte et al., 2023; He et al., 2022b; Rajput et al., 2024; Yuan et al., 2023; Mysore et al., 2023; Li et al., 2023c). For instance: LLM2-BERT4Rec (Harte et al., 2023) initializes BERT4Rec (Section 2.1.1) item embeddings using OpenAI embeddings of item texts; Query-SeqRec (He et al., 2022b) includes LLM query embeddings as inputs to a transformer-based recommender; and TIGER (Rajput et al., 2024) first uses an LLM to embed item text, then quantizes this embedding into a semantic ID, and finally trains a T5-based RS to generate new IDs given a user’s item ID history. Similarly, MINT (Mysore et al., 2023) and GPT4Rec (Li et al., 2023c) produce inputs for a dense retriever by prompting an LLM to generate a query given a user’s interaction history.
3.5. Conversational Recommendation
The recent advances in LLMs have made fully NL system-user dialogues a feasible and novel recommendation interface, bringing in a new stage of conversational recommendation (ConvRec) research. This direction studies the application of LLMs in multi-turn, multi-task, and mixed-initiative NL recommendation conversations (Friedman et al., 2023; Jannach et al., 2020). It introduces dialogue history as a rich new form of interaction data. Specifically, ConvRec includes the study and integration of diverse conversational elements such as dialogue management, recommendation, explanation, QA, critiquing, and preference elicitation (Jannach et al., 2020; Lyu et al., 2021). While some research (He et al., 2023) approaches ConvRec with a monolithic LLM such as GPT4, other works rely on an LLM to facilitate NL dialogue and integrate calls to a recommender module which generates item recommendations based on dialogue or interaction history (Li et al., 2018a; Chen et al., 2019b; Wang et al., 2022b; Yang et al., 2022). Further research makes progress towards ConvRec system architectures consisting of multiple tool-augmented LLM modules, including components for dialogue management, explanation generation, and retrieval (Friedman et al., 2023; Gao et al., 2023; Wang et al., 2023a, c).
4. Generative Multimodal Recommendation Systems
In recent years, users have come to expect richer levels of interactions than simple text or image queries, such as providing a picture of their desired product along with a natural language modification (e.g., a dress like the one in the picture but in red). Moreover, users want to better understand the recommendations they receive by visualizing how the product fits their use case (e.g., with a representation of how a garment might look on them, or how a furniture item might look in their room). Such levels of interaction require new RS that can discover the unique attributes or aspects hidden in each modality.
In this section, we discuss RS that utilize multiple data modalities. In Sections 4.1-4.2 we discuss motivations and challenges to the design of multimodal RS. In Sections 4.3-4.4 we review contrastive and generative approaches to multimodal RS, respectively.
4.1. Why Multimodal Recommendation?
Retailers often have multimodal information about their customers and products, including product descriptions, images and videos, customer reviews and purchase history. However, existing RS typically process each source independently and then combine the results by fusing unimodal relevance scores.
In practice, there are many use cases in which such a “late fusion” approach may be insufficient to satisfy the customer needs. One such use case is the cold start problem: when user behavioral data cannot be used to recommend existing products to new customers, or new products to existing customers, it is useful to gather diverse information about the items so that preference information can be transferred from existing products or customers to new ones.
Another use case occurs when different modalities are needed to understand the user request. For example, to answer the request “best metal and glass black coffee table under \$300 for my living room”, the system would need to reason about the appearance and shape of the item in context with the appearance and shape of other objects in the customer room, which cannot be achieved by searching with either text or image independently. Other examples of multimodal requests include an image or audio of the desired item together with text modification instructions (e.g., a song like the sound clip provided but in acoustic), or a complementary related product (e.g., a kickstand for the bicycle in the picture).
A third use case where multimodal understanding becomes crucial is when considering RS with complex outputs, like those featuring virtual try-on capabilities, or intelligent multimodal conversational shopping assistants.
4.2. Challenges to Multimodal Recommendation
The development of multimodal RS faces several challenges. First, collecting data to train multimodal systems (e.g., image-text-image triplets) is significantly harder than for unimodal systems. [^3] For example, data augmentation techniques used to define positive pairs for contrastive learning in the unimodal case do not apply to the the multimodal case. As a result, annotations for some modalities may be incomplete (Rahate et al., 2022). Second, combining different data modalities to improve recommendation results is not simple. For instance, existing contrastive learning approaches (Radford et al., 2021; Jia et al., 2021; Li et al., 2021, 2022a) map each data modality to a common latent space in which all modalities are approximately aligned. However, such approaches often capture information that is shared across modalities (e.g., text describing visual attributes), but they overlook complementary aspects that could benefit recommendations (e.g., text describing non visual attributes) (Guo et al., 2019). In general we would like the modalities to compensate for one another and result in a more complete joint representation. While fusion-based approaches (Li et al., 2021, 2022a) do learn a joint multimodal representation, ensuring the alignment of information that is shared and leaving some flexibility to capture complementary information across modalities remains a challenge. Third, learning multimodal models requires orders of magnitude more data than learning models for individual data modalities.
Despite these challenges, we believe multimodal generative models will become the standard approach. Indeed, recent literature shows significant advances on the necessary components to achieve effective multimodal generative models for RS, including (1) the use of LLMs and diffusion models to generate synthetic data for labeling purposes (Brooks et al., 2023; Rosenbaum et al., 2022; Nguyen et al., 2024), (2) high quality unimodal encoders and decoders (He et al., 2022a; Kirillov et al., 2023), (3) better techniques for aligning the latent spaces from multiple modalities into a shared one (Radford et al., 2021; Li et al., 2022a; Girdhar et al., 2023), (4) efficient re-parametrizations and training algorithms (Jang et al., 2016), and (5) techniques to inject structure to the learned latent space to make the problem tractable (Sohl-Dickstein et al., 2015).
4.3. Contrastive Multimodal Recommendation
As discussed before 4.2, learning multimodal generative models is very difficult because we need to not only learn a latent representation for each modality but also ensure that they are aligned. One way to address this challenge is to first learn an alignment between multiple modalities and then learn a generative model on “well-aligned” representations. In this subsection, we discuss two representative contrastive learning approaches: CLIP and ALBEF.
Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) is a popular approach, in which the task is to project images and associated text into the same point of the embedding space with parallel image and text encoders. This is achieved with a symmetric cross-entropy loss over the rows and columns of the cosine similarity matrix between all possible pairs of images and text in a training minibatch.
Align Before you Fuse (ALBEF) (Li et al., 2021) augments CLIP with a multimodal encoder that fuses the text and image embeddings, and proposes three objectives to pre-train the model: Image-text contrastive learning (ITC), masked language modeling (MLM), and image-text matching (ITM). The authors also introduce momentum distillation to provide pseudo-labels in order to compensate for the potentially incomplete or wrong text descriptions in the noisy web training data. Using their proposed architecture and training objectives, ALBEF obtains better results than CLIP in several zero-shot and fine-tuned multimodal benchmarks, despite using orders of magnitude less images for pre-training.
Contrastive-based alignment has shown impressive zero-shot classification and retrieval results (Novack et al., 2023; Baldrati et al., 2023; Hendriksen et al., 2022), and has been successfully fine-tuned to a multitude of tasks, such as object detection (Gu et al., 2021), segmentation (Zhou et al., 2023) or action recognition (Huang et al., 2024b). The same alignment objective has also been used between other modalities (Cheng et al., 2020; Hager et al., 2023; Huang, 2023), and with multiple modalities at the same time (Girdhar et al., 2023).
4.4. Generative Multimodal Recommendation
Despite their advantages, the performance of purely contrastive RS often suffers from data sparsity and uncertainty (Wang et al., 2022a). Generative models address these issues by imposing suitable structures on their latent spaces. Moreover, generative models allow for more complex recommendations, e.g., those requiring to synthesize an image. In what follows, we discuss thee representative generative approaches: VAEs, diffusion models, and multimodal LLMs.
Multimodal VAEs: While VAEs (see Section 2.1.2) could be applied directly to multimodal data, a better approach that leverages modality specific encoders and decoders trained on large corpus of data is to partition both the input and latent spaces per modality, say image and text. However, this approach reduces the multimodal VAE to two independent VAEs, one per modality. In ContrastVAE (Wang et al., 2022a), both modalities are aligned by adding a contrastive loss between the unimodal latent representations to the ELBO objective. Experiments show that ContrastVAE improves upon purely contrastive models by adequately modeling data uncertainty and sparsity, and being robust to perturbations in the latent space.
Diffusion models, explained in Section 2.4, are state-of-the-art models for image generation. While they can also be used for text generation, e.g., by using a discrete latent space with categorical transition probabilities (Austin et al., 2021), text encoders based on transformers or other sequence-to-sequence models are preferred in practice. As a consequence, multimodal models for both text and images, such as text-to-image generation models, combine text encoders with diffusion models for images. For instance, DALL-E (Ramesh et al., 2022) uses the CLIP embedding space as a starting point to generate novel images, and Stable Diffusion (Rombach et al., 2022) uses a UNet autoencoder separately pre-trained on a perceptual loss and a patch-based adversarial objective. Several works have built on and expanded diffusion models by increasing controllability of the generated results (Zhang et al., 2023c), consistency on the generated subjects identity (Ruiz et al., 2023), or for virtual try on applications (Zhu et al., 2023).
Multimodal LLMs (MLLM) provide a natural language interface for users to express their queries in multiple modalities, or even see responses in different modalities to help visualize the products. Given the complexity of training large generative models end-to-end, researchers typically assemble systems composed of discriminatively pre-trained encoders and decoders, usually connected by adaptation layers to ensure that unimodal representations are aligned. Another approach that involves little or no training is to allow a ”controller” LLM to use external foundation models, or tools, to deal with the multimodal input and output (Zhang et al., 2024). Then, instruction tuning is an important step to make LLMs useful task solvers. Llava (Liu et al., 2024) is a multimodal LLM that accepts both text and image inputs, and produces useful textual responses. The authors connect a CLIP encoder with an LLM decoder using a simple linear adaptation layer. In (Liu et al., 2023b) the authors change the connection layer from a linear projection to a two-layer MLP and obtain better results. Although MLLM research is still in its inception, some works already start using them in recommendation applications (Karra and Tulabandhula, 2024).
5. Evaluating for Impact and Harm
Evaluating RS is a complex and multifaceted task that goes beyond simply measuring a few key metrics of a single model. These systems are composed of one or more recommender models and various other ML and non-ML components, making it highly non-trivial to assess and evaluate the performance of an individual model. Moreover, these systems can have far-reaching impacts on users’ experiences, opinions, and actions, which may be difficult to quantify or predict, which adds to the challenge. The introduction of Gen-RecSys further complicates the evaluation process due to the lack of well-established benchmarks and the open-ended nature of their tasks. When evaluating RS, it is crucial to distinguish between two main targets of evaluation: the system’s performance and capabilities, and its potential for causing safety issues and societal harm. We review these targets, discuss evaluation metrics, and conclude with open challenges and future research directions.
5.1. Evaluating for Offline Impact
The typical approach to evaluating a model involves understanding its accuracy and efficiency in an offline setting, followed by live experiments.
5.1.1. Accuracy Metrics
The usual metrics used for discriminative tasks are recall@k, precision@k, NDCG@k, AUC, ROC, RMSE, MAE, etc. Many recent works on generative recommender models (e.g., (Kang et al., 2017, 2023; Rajput et al., 2024; Bao et al., 2023; He et al., 2023) incorporate such metrics for discriminative tasks.
For the generative tasks, we can borrow techniques from NLP. For example, the BLEU score is widely used for machine translation and can be useful for evaluating explanations(Geng et al., 2022), review generation, and conversational recommendations. The ROUGE score, commonly used for evaluating machine-generated summarization, could be helpful again for explanations or review summarization. Similarly, perplexity is another metric that could be broadly useful, including during the training process to ensure that the model is learning the language modeling component appropriately (Lu et al., 2021).
5.1.2. Computational Efficiency
Evaluating computational
efficiency is crucial for generative recommender models, both for training and inference, owing to their computational burden. This is an upcoming area of research.
5.1.3. Benchmarks
Many existing benchmark datasets popular in discriminative recommender models, such as Movielens (Harper and Konstan, 2015), Amazon Reviews (He and McAuley, 2016), Yelp Challenge(yel, 2014), Last.fm (Schedl, 2016), and Book-Crossing(Ziegler et al., 2005), are still useful in generative recommender models, but only narrowly. Some recent ones, like ReDial(Li et al., 2018b) and INSPIRED(Hayati et al., 2020), are useful datasets for conversational recommendations. (Deldjoo and Di Noia, 2024; Zhang et al., 2023a) propose benchmarks called cFairLLM and FaiRLLM, to evaluate consumer fairness in LLMs based on the sensitivity of pretrained LLMs to protected attributes in tailoring recommendations. We note that some benchmarks such as BigBench(bench authors, 2023) which are commonly used by the LLM community, have recommendations tasks. We believe it will be specifically useful for the recommendation community to develop new benchmarks for tasks unlocked by generative recommender models.
5.2. Online and Longitudinal Evaluations
Offline experiments may not capture an accurate picture because of the interdependence of the different models used in the system and other factors. So, A/B experiments help understand the model’s performance along several axes in real-world settings. Note that (Wang et al., 2023e) proposes a new paradigm of using simulation using agents to evaluate recommender models.
In addition to the short-term impact on engagement and satisfaction, the platform owners will be interested in understanding the long-term impact. This can be measured using business metrics such as revenue and engagement (time spent, conversions, etc.). At the same time, several metrics could be used to capture the impact on users (daily/monthly active users, user sentiment, safety, harm, etc.).
5.3. Conversational Evaluation
BLEU and perplexity are helpful metrics for conversational evaluation and need to be supplanted by task-specific metrics (such as recall) or objective-specific metrics (such as measuring diversity in responses(Li et al., 2015)). Strong LLMs could act as judges, but usually human evaluation is the ultimate choice here. Toolkits such as CRSLab(Kun Zhou, 2021) make it easy to build and evaluate conversational models. However, in many industrial use cases, the lack of labeled data makes this process challenging. Some works  (Friedman et al., 2023) use LLM-powered user simulations to generate data.
5.4. Evaluating for Harm
Milano et al. (2020) proposes six categories of harms associated with RS: content; privacy violations and data misuse; threats to human autonomy and well-being; transparency and accountability; harmful social effects such as filter bubbles, polarisation, and manipulability; and fairness. In addition to these, RS based on generative models may present new challenges (Weidinger et al., 2022; Bird et al., 2023):
  • LLMs use out-of-domain knowledge, introducing different sources of bias that are not easily captured by existing evaluation techniques (Shen et al., 2023; Deldjoo, 2024).
  • Heightened environmental impacts (Berthelot et al., 2024).
  • The automation of content creation and curation may displace human workers in industries such as journalism (De Cremer et al., 2023), creative writing, and content moderation, leading to social and economic disruption (Arguedas and Simon, 2023).
5.5. Holistic Evaluations
As mentioned above, thoroughly evaluating RS for offline metrics, online performance, and harm is highly non-trivial. Moreover, different stakeholders (e.g. platform owners and users)(Abdollahpouri et al., 2020; Sürer et al., 2018; Milano et al., 2021) may approach evaluation differently.
Given the complexity of evaluating Gen-RecSys, there is a significant opportunity for additional research and the development of specialized tooling. Moreover, it would benefit the community to draw inspiration from the HELM benchmark (Liang et al., 2022) and design a similar comprehensive evaluation framework tailored explicitly for Gen-RecSys.
6. Conclusions and Future Directions
While many directions for future work have been highlighted above, the following topics constitute especially important challenges and opportunities for Gen-RecSys:
  • RAG (cf. Section 3.3), including: data fusion for multiple (potentially subjective) sources such as reviews (Yu et al., 2021; Ye et al., 2023), end-to-end retriever-generator training (Lewis et al., 2020; Izacard and Grave, 2020; Borgeaud et al., 2022), and systematic studies of generative reranking alternatives (Qin et al., 2023).
  • Tool-augmented LLMs for conversational recommendation, focusing on architecture design for LLM-driven control of dialogue, recommender modules, external reasoners, retrievers, and other tools (Friedman et al., 2023; Wang et al., 2023a; Mialon et al., 2023; Bubeck et al., 2023), especially methods for proactive conversational recommendation.
  • Personalized Content Generation such as virtual try-on experiences (Zhu et al., 2023), which can allow users to visualize themselves wearing recommended clothing or accessories, improving customer satisfaction and reducing returns.
  • Red-teaming – in addition to the standard evaluations, real-world generative RS will have to undergo red-teaming (i.e., adversarial attacks) (Deng et al., 2023; Wang et al., 2023f; shu et al., 2024) before deployment to stress test the system for prompt injections, robustness, alignment verification, and other factors.
Despite being a short survey, in this work, we have attempted to provide a deeper understanding and exploration of the rich landscape of generative models within the field of recommendation systems. This survey extends the discussion beyond large language models (LLMs) to a broad spectrum of generative models and elucidates the nuanced intricacies of their applications across user-item interactions, textual data, and multimodal contexts. Additionally, it illuminates key evaluation challenges and dimensions, addressing the complexities of evaluating Gen-RecSys with considerations for performance, fairness, privacy, and societal impact, thereby establishing a new benchmark for future research in the domain.